# A Game-Theoretic Unital Classification Theorem for $C^{*}$-Algebras

### v1

@@ 2.3. Metric dynamic Ehrenfeucht-Fraïssé games
  • $r_{0}^{\Omega}(Aa,Bb):=\sup_{\varphi}|\varphi^{A}(a)-\varphi^{B}(b)|$ where $\varphi$ ranges over finitary quantifier-free $\Omega$-formulas;
  • $r_{\alpha+1}^{\Omega}(Aa,Bb):=\sup_{c\in A,d\in B}\inf_{c^{\prime}\in A,d^{\prime}\in B}(\max(r^{\Omega}_{\alpha}(Aac,Bbd^{\prime}),r^{\Omega}_{\alpha}(Aac^{\prime},Bbd)));$
@@
    $r^{\Omega}_{\gamma}(Aa,Bb):=\sup_{\beta<\gamma}r_{\beta}^{\Omega}(Aa,Bb).$
@@
[…]
Fact 2.11 ([yaacov2017metric, Proposition 4.3]).
@@
Fact 2.12 ([yaacov2017metric, Theorem 3.5]).
@@
While back-and-forth pseudo-distance only considers formulas respecting some fixed weak modulus $\Omega$, this restriction is overcome in [yaacov2017metric, Chapter 5]. Namely, the authors identify a so-called universal modulus $\Omega$ such that for any $L_{\omega_{1},\omega}$-sentence $\sigma$, there is an equivalent $\Omega$-sentence $\sigma^{\prime}$. The following fact summarizes this.
Fact 2.13 ([yaacov2017metric, Section 5]).
There exists a weak modulus $\Omega$ such that for any $L$-structures AA and $B$ with $A\equiv_{\alpha}^{bf,\Omega}B$ for each $\alpha<\omega_{1}$, it follows that $A\cong B$.
@@
We now characterize the pseudo-distance $r_{n,\alpha}$ in terms of a game. Denote by $\operatorname{LO}\subseteq\mathcal{P}(\omega)\times\mathcal{P}(\omega^{2})$ the set of linear orders on subsets of $\omega$. For $L$ a metric language, $L$-structures AA and $B$ and a countable linear order $(\lambda,\prec)\in\operatorname{LO}$ the metric dynamic Ehrenfeucht-Fraïssé Game $\operatorname{EFD}_{\prec,\varepsilon}(A,B)$ is played between Players I and II in the following way. The game clock is set first to $\beta_{0}:=\infty$. For $k\geq 1$, the $k$-th round is played as follows.
@@
    Player I plays either a pair $(\beta_{k},a_{k})$ or $(\beta_{k},b_{k})$ where $\beta_{k}\in\lambda$ with $\beta_{k}\prec\beta_{k-1}$, and $a_{k}\in A$ or $b_{k}\in B$.
@@
    Player II responds by playing $b_{k}\in B$ if I played an element in $A$, or by playing $a_{k}\in A$ if I played an element in $B$.
We call the tuple $(\beta_{k},(a_{1},\dots,a_{k}),(b_{1},\dots,b_{k}))$ the position at $k$. The game ends after round $k$ if $\beta_{k}$ is the minimum of $(\lambda,\prec)$ (if a minimum exists). In this case, Player II wins if and only if $r_{0,k}(Aa_{1}\dots a_{k},Bb_{1}\dots b_{k})<\varepsilon$. Note however that if $\prec$ is not well-ordered, the game may go on for (countably) infinitely many rounds, after which infinite tuples $a\in A^{\omega}$ and $b\in B^{\omega}$ will have been played. In this case, Player II has won if and only if $r_{0,k}(Aa_{1}\dots a_{k},Bb_{1}\dots b_{k})<\varepsilon$ for all $k<\omega$.
Note that while we have defined $\operatorname{EFD}_{\prec,\varepsilon}(A,B)$ for arbitrary $L$-structures, we will be interested only in separable ones, viewed as elements of the space $\operatorname{Mod}_{\omega}(L)$ defined in Section 2. In a game on separable structures $A,B\in\operatorname{Mod}_{\omega}(L)$, we require all moves $a_{k}$ and $b_{k}$ to be elements of the fixed countable dense subset of $A$ and $B$ respectively, which we identify with $a_{k},b_{k}\in\omega$. We will define what a winning strategy for Player II is in terms of separable structures, but this can of course be generalized to arbitrary $A$ and $B$. Intuitively, a strategy for II gives for each potential position at least one legal move that II can respond with. Such a strategy is winning for II if II is guaranteed to win as long as they play by this strategy.
@@
  • for all $\beta_{n+1}\prec\beta_{n}$ and $c\in\omega$, there are $d,d^{\prime}\in\omega$ such that $(\beta\beta_{n+1},ac,bd)\in S_{n+1}$ and $(\beta\beta_{n+1},ad^{\prime},bc)\in S_{n+1}$;
  • if $(\beta,a,b)\in S_{n}$, then $r_{0,n}(Aa,Bb)<\varepsilon$.
We write $\mathrm{II}\uparrow\operatorname{EFD}_{\prec,\varepsilon}(A,B)$ if there exists a winning strategy for Player $\mathrm{II}$, and denote the set of winning strategies for $\mathrm{II}$ by $\mathcal{S}_{\prec,\varepsilon}^{A,B}(\mathrm{II})\subseteq\prod_{n}\mathcal{P}(\omega^{n}\times\omega^{n}\times\omega^{n})$.
Analogously, we define the notion of a winning strategy for Player I. For $(S_{n})_{n}$ a strategy and $(\beta,a,b)\in S_{n}$, we write $S_{n+1}(\beta,a,b):=\{(\beta\beta_{n+1},ac,bd)\in S_{n+1}\}$ for the set of positions allowed by the strategy as a response to position $(\beta,a,b)$.
@@
Let $A,B\in\operatorname{Mod}_{\omega}(L)$ and $(\lambda,\prec)\in\operatorname{LO}$ and $\varepsilon>0$. A winning strategy for $\mathrm{I}$ in $\operatorname{EFD}_{\prec,\varepsilon}(A,B)$ is a sequence $(S_{n})_{n\geq 1}$ with $S_{n}\subseteq\omega^{n}\times\omega^{n}\times\omega^{n}$ such that
$\exists\beta\in\lambda\ \exists c\in\omega~\colon\big(\forall d\in\omega~\colon(\beta,c,d)\in S_{1}\vee(\beta,d,c)\in S_{1}\big)$
@@
  • If $\beta_{n}\neq\min(\lambda,\prec)$ then there exists $\beta_{n+1}\prec\beta_{n}$ and $c\in\omega$ such that either
    $S_{n+1}(\beta,a,b)=\{(\beta\beta_{n+1},ac,bd)\colon d\in\omega\}$
    or
    $S_{n+1}(\beta,a,b)=\{(\beta\beta_{n+1},ad^{\prime},bc)\colon d^{\prime}\in\omega\};$
  • If $\beta=\min_{\prec}(\omega)$, then $r_{0,n}(Aa,Bb)\geq\varepsilon$;
  • There exists $k<\omega$ and $(\beta\beta^{\prime},ac,bd)\in S_{n+k}$ such that $r_{0,n+k}(Aac,Bbd)\geq\varepsilon$.
@@
For $a\in\omega^{n}$ and $A\in\operatorname{Mod}_{\omega}(L)$, denote by $Aa$ the expansion of $A$ by an $n$-ary constant symbol interpreted as $a$. This is a separable structure in the language $L\cup\{c_{1},\dots,c_{n}\}$. Note that the game $\operatorname{EFD}_{\prec,\varepsilon}(Aa,Bb)$ is essentially equivalent to the game $\operatorname{EFD}_{n+\prec,\varepsilon}(A,B)$, with $a$ and $b$ as the forced first $n$ moves, since a winning strategy for II in either one of these games translates to one in the other. The metric dynamic Ehrenfeucht-Fraïssé game is defined precisely in a way that models the back-and-forth pseudo-metric. This is formalized in the following.
@@
We use induction on $\alpha$. Clearly, the statement holds for $\alpha=0$. For the successor step, note that
$r_{\alpha+1,n+1}(Aa,Bb)<\varepsilon\iff\forall c<\omega\ \exists d,d^{\prime}<\omega\colon(\max(r_{\alpha,n+1}^{A,B}(ac,bd),r_{\alpha,n+1}(Aad^{\prime},Bbc))<\varepsilon)$
@@
Suppose $(S_{n})$ is a winning strategy for II in $\operatorname{EFD}_{\alpha+1,\varepsilon}(Aa,Bb)$. Let $c\in\omega$. Then there is some $d\in\omega$ such that $(\alpha,c,d)\in S_{1}$. Now,
@@
defines a winning strategy for II in $\operatorname{EFD}_{\alpha,\varepsilon}(Aac,Bbd)$. The case for defining a winning strategy for II in $\operatorname{EFD}_{\alpha,\varepsilon}(Aad^{\prime},Bbc)$ is similar.
On the other hand, suppose for each $c<\omega$, there exists $d(c),d^{\prime}(c)<\omega$ and winning strategies $(S_{c,n})_{n}\in\mathcal{S}_{\alpha,\varepsilon}^{Aac,Bbd}(\mathrm{II})$ and $(S_{c,n}^{\prime})_{n}\in\mathcal{S}_{\alpha,\varepsilon}^{Aad^{\prime},Bbc}(\mathrm{II})$. Then $S_{1}:=\{(\beta,c,d(c))\colon c\in\omega,\beta\in\alpha\}\cup\{(\beta,d^{\prime}(c),c)\colon c\in\omega,\beta\in\alpha\}$ and
@@
witnesses $\mathrm{II}\uparrow\operatorname{EFD}_{\alpha+1}^{\varepsilon}(Aa,Bb)$.
For the limit step, it is enough to notice that whenever $\gamma<\omega_{1}$ is a limit ordinal, $\mathrm{II}\uparrow\operatorname{EFD}_{\gamma,\varepsilon}(Aa,Bb)$ if and only if $\mathrm{II}\uparrow\operatorname{EFD}_{\alpha,\varepsilon}(Aa,Bb)$ for all $\alpha<\gamma$. ∎
@@
For $A,B\in\operatorname{Mod}_{\omega}(L)$, $A\equiv_{\alpha}^{bf}B$ if and only if $\mathrm{II}$ has a winning strategy in $\operatorname{EFD}_{\alpha}(A,B)$.
@@
By definition, $A\equiv_{\alpha}^{bf}B$ if any only if $r_{0,\alpha}(A,B)=0$. The statement then follows immediately from Proposition 2.16 and the fact that a winning strategy for II in $\operatorname{EFD}_{\alpha}(A,B)$ has to contain a winning strategy in $\operatorname{EFD}_{\alpha,\varepsilon}(A,B)$ for every $\varepsilon>0$. ∎
@@
For any $\varepsilon>0$, $(\lambda,\prec)\in\operatorname{LO}$ and $A,B\in\operatorname{Mod}_{\omega}(L)$, the sets $\mathcal{S}_{\prec,\varepsilon}^{A,B}(\mathrm{I})$ and $\mathcal{S}_{\prec,\varepsilon}^{A,B}(\mathrm{II})$ are Borel in $\prod_{n}\mathcal{P}(\lambda^{n}\times\omega^{n}\times\omega^{n})$.
@@
This is immediate from the way the definitions above are written, together with the fact that the maps $r_{0,n}(A,\cdot,B,\cdot)\colon\omega^{n}\times\omega^{n}\to\mathbb{R}^{+}$ are Borel by Fact 2.11. ∎
We show now that the games we have defined are determined, that is, one of the two players always has a winning strategy (but not both).
@@
This follows from the fact that by definition, $\operatorname{EFD}_{\prec,\varepsilon(A,B)}$ is a closed game. That is, even though a gameplay can have infinitely many rounds, if I wins in a gameplay, then this is decided after a finite number rounds. It is a classical fact from descriptive set theory that such games are determined, see [MR1321597, Chapter II.20]. ∎
@@
$R_{\varepsilon}:=\{(A,B,\prec)\in\operatorname{Mod}_{\omega}(L)^{2}\times\operatorname{LO}\,\colon\,\mathrm{II}\uparrow\operatorname{EFD}_{\prec,\varepsilon}(A,B)\}$
@@
$Q_{\varepsilon}:=\{(A,B,\prec)\in\operatorname{Mod}_{\omega}(L)^{2}\times\operatorname{LO}\,\colon\,\mathrm{I}\uparrow\operatorname{EFD}_{\prec,\varepsilon}(A,B)\}.$
@@
[…]
@@
We conclude with Theorem D, noting that any countable ordinal $\alpha<\omega_{1}$ is isomorphic to some $\prec\in\operatorname{LO}$.
@@
Taking $\varepsilon$ to range over $\{n^{-1}\colon n<\omega\}$, $R$ is Borel as a countable intersection of Borel sets by 2.20. The equivalence holds by Proposition 2.16. ∎

### v2

@@ 2.3. Metric dynamic Ehrenfeucht-Fraïssé games
  • $r_{0,n}^{\Omega}(Aa,Bb):=\sup_{\varphi}|\varphi^{A}(a)-\varphi^{B}(b)|$ where $\varphi$ ranges over finitary quantifier-free $\Omega$-formulas;
  • $r_{\alpha+1,n}^{\Omega}(Aa,Bb):=\sup_{c\in A,d\in B}\inf_{c^{\prime}\in A,d^{\prime}\in B}(\max(r^{\Omega}_{\alpha,n+1}(Aac,Bbd^{\prime}),r^{\Omega}_{\alpha,n+1}(Aac^{\prime},Bbd)));$
@@
    $r^{\Omega}_{\gamma,n}(Aa,Bb):=\sup_{\beta<\gamma}r_{\beta,n}^{\Omega}(Aa,Bb).$
@@
We restrict attention to separable $L$-structures, which we view as elements of the Borel space $\operatorname{Mod}_{\omega}(L)$, as discussed above Definition 2.7. In particular, if $A,B\in\operatorname{Mod}_{\omega}(L)$ and $a,b\in\omega^{n}$, we write $r_{\alpha,n}^{\Omega}(Aa,Bb)$ for $r_{\alpha,n}^{\Omega}(\widehat{A}a,\widehat{B}b)$. By continuity of $r_{\alpha,n}^{\Omega}$ [2, Lemma 3.2], it is sufficient in the separable case to restrict the range of supremum and infimum in the successor step to a distinguished countable dense set. An inductive argument shows the following.
Fact 2.11 ([2, Proposition 4.3]).
@@
Fact 2.12 ([2, Theorem 3.5]).
@@
While back-and-forth pseudo-distance only considers formulas respecting some fixed weak modulus $\Omega$, this restriction is overcome in [2, Chapter 5]. Namely, the authors prove existence of a so-called universal modulus $\Omega$ such that for any $L_{\omega_{1},\omega}$-sentence $\sigma$, there is an equivalent $L_{\omega_{1},\omega}$-$\Omega$-sentence $\sigma^{\prime}$. The following fact summarizes this.
Fact 2.13 ([2, Section 5]).
There exists a weak modulus $\Omega$ such that for any $L$-structures AA and $B$,
$\big(\forall\alpha<\omega_{1}:~A\equiv_{\alpha}^{bf,\Omega}B\big)\iff A\cong B.$
@@
We now turn to characterizing the pseudo-distance $r_{n,\alpha}$ in terms of a game. Denote by $\operatorname{LO}\subseteq\mathcal{P}(\omega)\times\mathcal{P}(\omega^{2})$ the set of linear orders on subsets of $\omega$. For $L$ a metric language, $L$-structures AA and $B$ and $(\lambda,\prec)\in\operatorname{LO}$ the metric dynamic Ehrenfeucht-Fraïssé Game $\operatorname{EFD}_{\prec,\varepsilon}(A,B)$ is played between Players I and II in the following way. The game clock is set first to $\beta_{0}:=\infty$. For $k\geq 1$, the $k$-th round is played as follows.
@@
    Player I plays either a pair $(\beta_{k},a_{k})$ or a pair $(\beta_{k},b_{k})$, where $\beta_{k}\in\lambda$ with $\beta_{k}\prec\beta_{k-1}$, and $a_{k}\in A$ or $b_{k}\in B$.
@@
    Player II responds by playing $b_{k}\in B$ if I played an element $a_{k}\in A$, or by playing $a_{k}\in A$ if I played an element $b_{k}\in B$.
We call the tuple $(\beta_{k},(a_{1},\dots,a_{k}),(b_{1},\dots,b_{k}))$ the position at $k$. The game ends after round $k$ if $\beta_{k}$ is the minimum of $(\lambda,\prec)$ (if a minimum exists). In this case, Player II wins if and only if $r_{0,k}(Aa_{1}\dots a_{k},Bb_{1}\dots b_{k})<\varepsilon$. Note that if $\prec$ is not well-ordered, the game may go on for countably infinitely many rounds, after which infinite tuples $a\in A^{\omega}$ and $b\in B^{\omega}$ will have been played and the game ends. In this case, Player II wins if and only if $r_{0,k}(Aa_{1}\dots a_{k},Bb_{1}\dots b_{k})<\varepsilon$ for all $k<\omega$.
While we have defined $\operatorname{EFD}_{\prec,\varepsilon}(A,B)$ for arbitrary $L$-structures, we will be interested only in separable ones, viewed as elements of the space $\operatorname{Mod}_{\omega}(L)$ defined in Section 2. In a game on separable structures $A,B\in\operatorname{Mod}_{\omega}(L)$, we require all moves $a_{k}$ and $b_{k}$ to be elements of the fixed countable dense subset of $A$ and $B$ respectively, which we identify with $a_{k},b_{k}\in\omega$. We will define what a winning strategy for Player II is in terms of separable structures, but this can straight-forwardly be generalized to arbitrary $L$-structures. Intuitively, a strategy for II gives for each potential position of the game at least one legal move that II can respond with. Such a strategy is winning for II if II is guaranteed to win as long as they play by this strategy.
@@
  • $\forall\beta_{n+1}\prec\beta_{n}\ \forall c\in\omega\ \exists d,d^{\prime}\in\omega~\colon(\beta\beta_{n+1},ac,bd),(\beta\beta_{n+1},ad^{\prime},bc)\in S_{n+1}$;
  • $r_{0,n}(Aa,Bb)<\varepsilon$.
We write $\mathrm{II}\uparrow\operatorname{EFD}_{\prec,\varepsilon}(A,B)$ if there exists a winning strategy for Player $\mathrm{II}$, and we denote the set of winning strategies for $\mathrm{II}$ by $\mathcal{S}_{\prec,\varepsilon}^{A,B}(\mathrm{II})\subseteq\prod_{n}\mathcal{P}(\omega^{n}\times\omega^{n}\times\omega^{n})$.
Analogously, we define the notion of a winning strategy for Player I. For $(S_{n})_{n\in\mathbb{N}}\in\prod_{n\in\mathbb{N}}\omega^{n}\times\omega^{n}\times\omega^{n}$ and $(\beta,a,b)\in S_{n}$, we write $S_{n+1}(\beta,a,b):=\{(\beta\beta_{n+1},ac,bd)\in S_{n+1}\}$.
@@
Let $A,B\in\operatorname{Mod}_{\omega}(L)$, $(\lambda,\prec)\in\operatorname{LO}$ and $\varepsilon>0$. A winning strategy for $\mathrm{I}$ in $\operatorname{EFD}_{\prec,\varepsilon}(A,B)$ is a sequence $(S_{n})_{n\in\mathbb{N}}$ with $S_{n}\subseteq\omega^{n}\times\omega^{n}\times\omega^{n}$ such that
  • $\exists\beta\in\lambda\ \exists c\in\omega~\colon\big(\forall d\in\omega~\colon(\beta,c,d)\in S_{1}\big)\ \vee\ \big(\forall d\in\omega~\colon(\beta,d,c)\in S_{1}\big)$,
@@
  • if $\beta_{n}\neq\min(\lambda,\prec)$ then $\exists\beta_{n+1}\prec\beta_{n}\ \exists c\in\omega$ s.t. $S_{n+1}(\beta,a,b)=\{(\beta\beta_{n+1},ac,bd)\colon d\in\omega\}$ or $S_{n+1}(\beta,a,b)=\{(\beta\beta_{n+1},ad^{\prime},bc)\colon d^{\prime}\in\omega\};$
  • if $\beta_{n}=\min(\lambda,\prec)$ then $r_{0,n}(Aa,Bb)\geq\epsilon$;
  • $\exists k<\omega\ \exists(\beta\beta^{\prime},ac,bd)\in S_{n+k}~\colon\ r_{0,n+k}(Aac,Bbd)\geq\varepsilon$.
@@
For $a\in\omega^{n}$ and $A\in\operatorname{Mod}_{\omega}(L)$, denote by $Aa$ the expansion of $A$ by an $n$-ary constant symbol interpreted as $a$. This is a separable structure in the language $L\cup\{c_{1},\dots,c_{n}\}$. Note that the game $\operatorname{EFD}_{\prec,\varepsilon}(Aa,Bb)$ is equivalent to the game $\operatorname{EFD}_{n+\prec,\varepsilon}(A,B)$, with $a$ and $b$ as the forced first $n$ moves, in the sense that a winning strategy for II in either one of these games translates to one in the other.
The metric dynamic Ehrenfeucht-Fraïssé game is defined precisely in a way that models the back-and-forth pseudo-metric. This is formalized in the following.
@@
We use induction on $\alpha$. Clearly, the statement holds for $\alpha=0$.
For the limit step, it is enough to observe that whenever $\gamma<\omega_{1}$ is a limit ordinal, $\mathrm{II}\uparrow\operatorname{EFD}_{\gamma,\varepsilon}(Aa,Bb)$ if and only if $\mathrm{II}\uparrow\operatorname{EFD}_{\alpha,\varepsilon}(Aa,Bb)$ for all $\alpha<\gamma$.
For the successor step, note that
$r_{\alpha+1,n}(Aa,Bb)<\varepsilon\iff\forall c<\omega\ \exists d,d^{\prime}<\omega\ \colon\ \max\big(r_{\alpha,n+1}^{A,B}(ac,bd),r_{\alpha,n+1}(Aad^{\prime},Bbc)\big)<\varepsilon$
@@
Suppose $(S_{n})_{n\in\mathbb{N}}$ is a winning strategy for II in $\operatorname{EFD}_{\alpha+1,\varepsilon}(Aa,Bb)$. Let $c\in\omega$. Then there is some $d\in\omega$ such that $(\alpha,c,d)\in S_{1}$. Now, setting for each $n\in\mathbb{N}$,
@@
defines a winning strategy for II in $\operatorname{EFD}_{\alpha,\varepsilon}(Aac,Bbd)$. A winning strategy for II in $\operatorname{EFD}_{\alpha,\varepsilon}(Aad^{\prime},Bbc)$ can be found analogously.
On the other hand, suppose for each $c<\omega$, there exists $d(c),d^{\prime}(c)<\omega$ and winning strategies $(S_{c,n})_{n}\in\mathcal{S}_{\alpha,\varepsilon}^{Aac,Bbd}(\mathrm{II})$ and $(S_{c,n}^{\prime})_{n}\in\mathcal{S}_{\alpha,\varepsilon}^{Aad^{\prime},Bbc}(\mathrm{II})$. Then setting
$S_{1}:=\{(\beta,c,d(c))\colon c\in\omega,\beta\in\alpha\}\cup\{(\beta,d^{\prime}(c),c)\colon c\in\omega,\beta\in\alpha\}$
and
@@
yields a winning strategy for $\mathrm{II}$ in $\operatorname{EFD}_{\alpha+1}^{\varepsilon}(Aa,Bb)$. ∎
@@
For $A,B\in\operatorname{Mod}_{\omega}(L)$, $A\equiv_{\alpha}^{bf}B$ if and only if $\mathrm{II}\uparrow\operatorname{EFD}_{\alpha}(A,B)$.
@@
By definition, $A\equiv_{\alpha}^{bf}B$ if any only if $r_{0,\alpha}(A,B)=0$. The statement then follows immediately from Proposition 2.16 and the fact that a winning strategy for II in $\operatorname{EFD}_{\alpha}(A,B)$ is precisely a winning strategy for $\mathrm{II}$ in $\operatorname{EFD}_{\alpha,\varepsilon}(A,B)$ for every $\varepsilon>0$. ∎
@@
For any $\varepsilon>0$, $(\lambda,\prec)\in\operatorname{LO}$ and $A,B\in\operatorname{Mod}_{\omega}(L)$, the sets $\mathcal{S}_{\prec,\varepsilon}^{A,B}(\mathrm{I})$ and $\mathcal{S}_{\prec,\varepsilon}^{A,B}(\mathrm{II})$ are Borel in $\prod_{n\in\mathbb{N}}\mathcal{P}(\omega^{n}\times\omega^{n}\times\omega^{n})$.
@@
This is immediate from the presentation of Definition 2.14 and Definition 2.15, together with the fact that the maps $r_{0,n}(A,\cdot,B,\cdot)\colon\omega^{n}\times\omega^{n}\to\mathbb{R}^{+}$ are Borel by Fact 2.11. ∎
We show now that the metric dynamic EF-games are determined, that is, one of the two players always has a winning strategy (but not both).
@@
This follows from the fact that by definition, $\operatorname{EFD}_{\prec,\varepsilon(A,B)}$ is a closed game. That is, even though a gameplay can have infinitely many rounds, if I wins in a gameplay, then this is decided after a finite number rounds. It is a classical fact from descriptive set theory that such games are determined, see for example [39, Chapter II.20]. ∎
@@
$Q_{\varepsilon}:=\{(A,B,\prec)\in\operatorname{Mod}_{\omega}(L)^{2}\times\operatorname{LO}\,\colon\,\mathrm{I}\uparrow\operatorname{EFD}_{\prec,\varepsilon}(A,B)\}$
@@
$R_{\varepsilon}:=\{(A,B,\prec)\in\operatorname{Mod}_{\omega}(L)^{2}\times\operatorname{LO}\,\colon\,\mathrm{II}\uparrow\operatorname{EFD}_{\prec,\varepsilon}(A,B)\}.$
@@
$\big\{(A,B,(\lambda,\prec),(S_{n})_{n})\ \colon\ (S_{n})_{n}\in\mathcal{S}_{\prec,\varepsilon}^{A,B}(\mathrm{II})\big\}$
@@
We conclude with the following key result, amounting to Theorem D. Note that any countable ordinal $\alpha<\omega_{1}$ is isomorphic to some $(\lambda,\prec)\in\operatorname{LO}$.
@@
Taking $\varepsilon$ to range over $\{n^{-1}\colon n<\omega\}$, $R$ is Borel as a countable intersection of Borel sets by Lemma 2.20. The equivalence holds by Proposition 2.16. ∎